# OptiBench Meets ReSocratic: Measure and Improve LLMs for Optimization Modeling

### v3

@@ Abstract
Large language models (LLMs) have exhibited their problem-solving abilities in mathematical reasoning. Solving realistic optimization (OPT) problems in application scenarios requires advanced and applied mathematics ability. However, current OPT benchmarks that merely solve linear programming are far from complex realistic situations. In this work, we propose OptiBench, a benchmark for end-to-end optimization problem-solving with human-readable inputs and outputs. OptiBench contains rich optimization problems, including linear and nonlinear programming with or without tabular data, which can comprehensively evaluate LLMs’ solving ability. In our benchmark, LLMs are required to call a code solver to provide precise numerical answers. Furthermore, to alleviate the data scarcity for optimization problems, and to bridge the gap between open-source LLMs on a small scale (e.g., Llama-3-8b) and closed-source LLMs (e.g., GPT-4), we further propose a data synthesis method namely ReSocratic. Unlike general data synthesis methods that proceed from questions to answers, ReSocratic first incrementally synthesizes formatted optimization demonstrations with mathematical formulations step by step and then back-translates the generated demonstrations into questions. Based on this, we synthesize the ReSocratic-29k dataset. We further conduct supervised fine-tuning with ReSocratic-29k on multiple open-source models. Experimental results show that ReSocratic-29k significantly improves the performance of open-source models. The code and data can be found at https://github.com/yangzhch6/ReSocratic.
@@ 2 Related Work
Data Synthesis for Mathematical Reasoning. Improving the performance of language models in mathematical reasoning tasks significantly depends on increasing the quantity of fine-tuning data for LLMs. A substantial body of work has been dedicated to these areas (Yu et al., 2023; Li et al., 2023; Yuan et al., 2023; Yue et al., 2023; Zhao et al., 2024; Lu et al., 2024b; Liu & Yao, 2024). One notable approach is Rejection Sampling Fine-Tuning (RFT; Yuan et al. 2023), which employs supervised models to generate and collect correct reasoning paths, creating augmented fine-tuning datasets. MAmmoTH (Yue et al., 2023) utilizes RFT with GPT-4 to gather both Chain-of-Thought solutions in natural language and Program-of-Thought solutions in formal language. Similarly, MetaMath (Yu et al., 2023) focuses on data augmentation for both question and answer texts. MathGenie (Lu et al., 2024b) collects a vast amount of data through open-source language models. While there has been significant progress in synthesizing data for informal mathematical reasoning, efforts have also been made to address formal reasoning through the use of formal languages and compilers (Xiong et al., 2023; Huang et al., 2024b; Lu et al., 2024a). However, a major challenge remains: there is a scarcity of high-quality data for optimization problems. This lack of data limits the direct application of these prior approaches to optimization contexts.
@@
Socratic Method. The Socratic method [^1] https://en.wikipedia.org/wiki/Socratic_method is a critical thinking method with dialogic disassembled multi-step sub-questions and answers cultivating in answering a complex question. This method has been applied by current language model techniques for advanced reasoning tasks, such as prompting step-wise reasoning (Qi et al., 2023; Chang, 2023; Shridhar et al., 2022), multi-agent interaction (Zeng et al., 2023), and discovering math knowledge (Dong et al., 2023). For example, Qi et al. (2023) proposes a divide-and-conquer style algorithm that mimics recursive thinking by asking Socratic questions, it thus relieves the reliance on the initial decision as chain-of-thought (CoT) and achieves performance improvements on several complex reasoning tasks. Another line of work (Ang et al., 2023; Cobbe et al., 2021) applies the Socratic method for fine-grained dataset construction. GSM8K Socratic dataset [^2] https://github.com/openai/grade-school-math?tab=readme-ov-file#socratic-dataset (Cobbe et al., 2021) is the most related work to our paper. They inject automatically generated “Socratic sub-questions” before each step, resulting in fine-grained math data. To construct a step-by-step benchmark for optimization problem solving with intermediate solutions, in this work, we explore the Socratic method to synthesize optimization problems. Unlike the previous study, we propose a reverse Socratic approach (ReSocratic) that generates optimization problems from the answer back to a question, and we demonstrate its superiority to traditional forward Socratic synthesis.
@@ 3 OptiBench: Benchmark for Optimization Modeling
Data Collection and Annotation. In the data annotation stage, we assign workers to collect questions from textbooks (Bertsimas & Tsitsiklis, 1997; Conforti et al., 2014; Wolsey, 2020), and a university’s course assignments and examinations. The workers are required to collect the questions first and save the text of the question in markdown format. If the question contains a table, convert the table to markdown format as well. If there is a standard answer to the question, it is also saved; if not, we let the worker complete the answer in natural language form. We require our workers to write Python code, call the pyscipopt [^3] https://github.com/scipopt/PySCIPOpt solver to solve each problem, and ask them to output the values of the variables and optimization targets at the end of the code. These numerical solutions are recorded in the dataset as ground truth answers. Additionally, for each collected sample, we require two additional annotators to participate in determining whether the annotation is correct and to correct any samples with incorrect annotations.
Figure 2 shows a mixed integer nonlinear programming sample of OptiBench. For each sample, we provide the “Question” and “Results”. For every problem, we ask the workers to judge the uniqueness of the optimal solution. If the workers determine that the solution is unique, the “Results” will contain the description (colored in orange in Figure 2) and optimal value (colored in green in Figure 2) of all the variables and optimization objective; otherwise “Results” contains only the optimal value of the optimization objective.
AI4MATH Contest. A preliminary version of OptiBench is released in April 2024 as a contest track [^4] https://www.codabench.org/competitions/2438/ of ICML 2024 Challenge on Automated Math Reasoning [^5] https://sites.google.com/view/ai4mathworkshopicml2024/challenges.
@@ 4 ReSocratic: Data Synthesis with Reverse Socratic
2) New Demonstration Synthesis. The ReSocratic method prompts an LLM to generate new demonstrations based on the seeds. We sample 2 seeds each time from the pool to form the synthesis prompt as shown in Appendix D.2. The LLM will follow the given prompt to generate new demonstrations step by step. The generated data is rigorously selected via (1) each intermediate step should follow the format (as shown in the first step) by a rule filter and (2) the overall demonstration does not overlap with generated ones by a similarity filter. For all generated demonstrations, we set a similarity filter, which converts all texts into TF-IDF vectors and filters out demonstrations with cosine similarity higher than a threshold. The procedure is shown in Figure 3(b).
3) Question and Code Translation. We construct a code generation prompt to solve the mathematical formulations in the synthetic demonstrations and output the optimal solution results. If the code runs incorrectly, we delete this demonstration. Next, to acquire the questions in plain text format and the questions in table format, we construct two back-translation prompts. All the prompts are shown in Appendix D. Then, for each generated demonstration starting from the third step, we translate it into a question-code pair, as shown in Figure 3(b). Each generated code will be executed automatically to ensure there are no bugs.
@@ 5.1 Baselines and Settings
Evaluation Setting. The evaluation metric of our OptiBench is the answer accuracy, as detailed in Section 2. We show the solving accuracy of the four data types along with the code pass rate. We evaluate LLMs under three settings: Zero-shot, Few-shot, and Supervised Fine-Tuning (SFT) setting. We provide the zero-shot prompt and the few-shot prompt to solve the problem in Appendix D.1.
Baselines. We select GPT-3.5-Turbo (Brown et al., 2020), GPT-4 (Achiam et al., 2023), Llama-2-7B-Chat (Touvron et al., 2023b), Llama-3-8B-Instruct (Team, 2024) and Llama-3-70B-Instruct (Team, 2024) as the baselines of zero-shot and few-shot settings. For the SFT setting, we use Llama-2-7B-Chat and textbfLlama-3-8B-Instruct.
@@
Fine-tuning Setting. For a given language model, we utilize our contributed ReSocratic-29k to conduct supervised fine-tuning. Specifically, we construct the training sample as follows:
"user": "{Question}"
"assistant":{Code}"
We replace “{Question}” and “{Code}” with the synthetic question and the verified code in our ReSocratic-29k to form the training sample. Based on this, we conduct fine-tuning experiments on two A800 GPUs, the epoch is set as 3, the learning rate is $2e^{-5}$, and the batch size is 128.
@@ 5.3 Main Results
Llama-3-70B-Instruct 79.2% 57.5% 33.8% 32.0% 62.5% 91.2%
DeepSeek-V2 79.5% 56.3% 27.1% 32.0% 61.0% 85.5%
GPT-3.5-Turbo 75.4% 40.0% 28.6% 26.0% 56.4% 93.2%
GPT-4 80.7% 71.3% 34.6% 34.0% 65.5% 88.3%
@@ 5.4 Performance Analysis on Data Split
1) Tabular questions are harder. We find from the experimental results, that it is more difficult to solve table questions than no-table questions, and nonlinear questions are more difficult than linear questions.
2) Nonlinear problem is harder than linear problem. According to Table 2, for all language models, the performance of nonlinear questions is significantly lower than that of linear questions. In Figure 6 (a), we show the average performance of linear problems and nonlinear problems for GPT-4 and GPT-3.5-Turbo. The performance is 66.9% for linear problems and 30.8% for nonlinear problems.
@@ Appendix A More Comparisons with Other Benchmarks
Appendix B More Detail of ReSocratic

### v4

@@ Abstract
Large language models (LLMs) have exhibited their problem-solving abilities in mathematical reasoning. Solving realistic optimization (OPT) problems in application scenarios requires advanced and applied mathematics ability. However, current OPT benchmarks that merely solve linear programming are far from complex realistic situations. In this work, we propose OptiBench, a benchmark for end-to-end optimization problem-solving with human-readable inputs and outputs. OptiBench contains rich optimization problems, including linear and nonlinear programming with or without tabular data, which can comprehensively evaluate LLMs’ solving ability. In our benchmark, LLMs are required to call a code solver to provide precise numerical answers. Furthermore, to alleviate the data scarcity for optimization problems, and to bridge the gap between open-source LLMs on a small scale (e.g., Llama-3-8b) and closed-source LLMs (e.g., GPT-4), we further propose a data synthesis method namely ReSocratic. Unlike general data synthesis methods that proceed from questions to answers, ReSocratic first incrementally synthesizes formatted optimization demonstrations with mathematical formulations step by step and then back-translates the generated demonstrations into questions. Based on this, we synthesize the ReSocratic-29k dataset. We further conduct supervised fine-tuning with ReSocratic-29k on multiple open-source models. Experimental results show that ReSocratic-29k significantly improves the performance of open-source models.
@@ 2 Related Work
Data Synthesis for Mathematical Reasoning. Improving the performance of language models in mathematical reasoning tasks significantly depends on increasing the quantity of fine-tuning data for LLMs. A substantial body of work has been dedicated to these areas (Yu et al., 2023; Liu & Yao, 2024; Li et al., 2023; Yuan et al., 2023; Lu et al., 2024b; Yue et al., 2023). One notable approach is Rejection Sampling Fine-Tuning (RFT; Yuan et al. 2023), which employs supervised models to generate and collect correct reasoning paths, creating augmented fine-tuning datasets. MAmmoTH (Yue et al., 2023) utilizes RFT with GPT-4 to gather both Chain-of-Thought solutions in natural language and Program-of-Thought solutions in formal language. Similarly, MetaMath (Yu et al., 2023) focuses on data augmentation for both question and answer texts. MathGenie (Lu et al., 2024b) collects a vast amount of data through open-source language models. While there has been significant progress in synthesizing data for informal mathematical reasoning, efforts have also been made to address formal reasoning through the use of formal languages and compilers (Xiong et al., 2023; Huang et al., 2024b; Lu et al., 2024a). However, a major challenge remains: there is a scarcity of high-quality data for optimization problems. This lack of data limits the direct application of these prior approaches to optimization contexts.
@@
Socratic Method. The Socratic method (Gose, 2009; Scholle, 2020) is a critical thinking method with dialogic disassembled multi-step sub-questions and answers cultivating in answering a complex question. This method has been applied by current language model techniques for advanced reasoning tasks, such as prompting step-wise reasoning (Qi et al., 2023; Chang, 2023; Shridhar et al., 2022), multi-agent interaction (Zeng et al., 2023), and discovering math knowledge (Dong et al., 2023). For example, Qi et al. (2023) proposes a divide-and-conquer style algorithm that mimics recursive thinking by asking Socratic questions, it thus relieves the reliance on the initial decision as chain-of-thought (CoT) and achieves performance improvements on several complex reasoning tasks. Chen & Lampouras (2023); Xie et al. (2023) utilize back-translation as core modules in their data generation frameworks, but our ReSocratic places more emphasis on step-by-step reverse construction of the chain-of-thought from scratch, with back-translation being one tiny step in our framework. Another line of work (Ang et al., 2023; Cobbe et al., 2021) applies the Socratic method for fine-grained dataset construction. GSM8K Socratic dataset [^1] https://github.com/openai/grade-school-math?tab=readme-ov-file#socratic-dataset (Cobbe et al., 2021) is the most related work to our paper. They inject automatically generated “Socratic sub-questions” before each step, resulting in fine-grained math data. To construct a step-by-step benchmark for optimization problem solving with intermediate solutions, in this work, we explore the Socratic method to synthesize optimization problems. Unlike the previous study, we propose a reverse Socratic approach (ReSocratic) that generates optimization problems from the answer back to a question, and we demonstrate its superiority to traditional forward Socratic synthesis.
@@ 3 OptiBench: Benchmark for Optimization Modeling
Data Collection and Annotation. In the data annotation stage, we assign workers to collect questions from textbooks (Bertsimas & Tsitsiklis, 1997; Conforti et al., 2014; Wolsey, 2020), and a university’s course assignments and examinations. We require our workers to write Python code, call the pyscipopt [^2] https://github.com/scipopt/PySCIPOpt solver to solve each problem, and ask them to output the values of the variables and optimization targets at the end of the code. Figure 2 shows a mixed integer nonlinear programming sample of OptiBench. For each sample, we provide the “Question” and “Results”. More details of data collection and annotation are shown in Appendix D.
@@ 4 ReSocratic: Data Synthesis with Reverse Socratic
2) New Demonstration Synthesis. The ReSocratic method prompts an LLM to generate new demonstrations based on the seeds. We sample 2 seeds each time from the pool to form the synthesis prompt as shown in Appendix E.2. The LLM will follow the given prompt to generate new demonstrations step by step. The generated data is rigorously selected via (1) each intermediate step should follow the format (as shown in the first step) by a rule filter and (2) the overall demonstration does not overlap with generated ones by a similarity filter. For all generated demonstrations, we set a similarity filter, which converts all texts into TF-IDF vectors and filters out demonstrations with cosine similarity higher than a threshold. The procedure is shown in Figure 3(b).
3) Question and Code Translation. We construct a code generation prompt to solve the mathematical formulations in the synthetic demonstrations and output the optimal solution results. If the code runs incorrectly, we delete this demonstration. Next, to acquire the questions in plain text format and the questions in table format, we construct two back-translation prompts. All the prompts are shown in Appendix E. Then, for each generated demonstration starting from the third step, we translate it into a question-code pair, as shown in Figure 3(b). Each generated code will be executed automatically to ensure there are no bugs.
@@ 5.1 Baselines and Settings
Evaluation Setting. The evaluation metric of our OptiBench is the answer accuracy, as detailed in Section 2. We show the solving accuracy of the four data types along with the code pass rate. We evaluate LLMs under three settings: Zero-shot, Few-shot, and Supervised Fine-Tuning (SFT) setting. We provide the zero-shot prompt and the few-shot prompt to solve the problem in Appendix E.1.
Baselines. We select GPT-family (Brown et al., 2020; Achiam et al., 2023), Llama-family (Touvron et al., 2023b; Team, 2024), Qwen2 Yang et al. (2024), Mistral-v0.3 Jiang et al. (2023), and DeepSeek-family (DeepSeek-AI, 2024) as the baselines in zero-shot and few-shot settings. For the SFT setting, we use Llama-2-7B-Chat and textbfLlama-3-8B-Instruct.
@@
Fine-tuning Setting. For a given language model, we utilize our contributed ReSocratic-29k to conduct supervised fine-tuning. Based on this, we conduct fine-tuning experiments on two A800 GPUs, the epoch is set as 3, the learning rate is $2e^{-5}$, and the batch size is 128.
@@ 5.3 Main Results
Qwen2-7b-Instruct 65.5% 27.5% 18.8% 14.0% 46.0% 87.6%
DeepSeek-V2.5 79.5% 71.3% 40.6% 48.0% 67.3% 91.2%
GPT-4o-mini 74.6% 52.5% 14.3% 34.0% 55.0% 74.4%
GPT-4o 81.0% 63.8% 50.4% 50.0% 69.4% 91.7%
@@ 5.4 Performance Analysis on Data Split
1) Tabular questions are harder. It is more difficult to solve table questions than no-table questions, and nonlinear questions are more difficult than linear questions.
2) Nonlinear problem is harder. According to Table 2, for all language models, the performance of nonlinear questions is significantly lower than that of linear questions. In Figure 6 (a), we show the average performance of linear problems and nonlinear problems for GPT-4 and GPT-3.5-Turbo. The performance is 66.9% for linear problems and 30.8% for nonlinear problems.
@@ 6 Conclusion
Acknowledgments
Jing Tang’s work is partially supported by National Key R&D Program of China under Grant No. 2023YFF0725100 and No. 2024YFA1012701, by the National Natural Science Foundation of China (NSFC) under Grant No. 62402410 and No. U22B2060, by Guangdong Provincial Project (No. 2023QN10X025), by Guangdong Basic and Applied Basic Research Foundation under Grant No. 2023A1515110131, by Guangzhou Municipal Science and Technology Bureau under Grant No. 2023A03J0667 and No. 2024A04J4454, by Guangzhou Municipal Education Bureau (No. 2024312263), and by Guangzhou Municipality Big Data Intelligence Key Lab (No. 2023A03J0012), Guangzhou Industrial Information and Intelligent Key Laboratory Project (No. 2024A03J0628) and Guangzhou Municipal Key Laboratory of Financial Technology Cutting-Edge Research (No. 2024A03J0630).
@@ Appendix A More Comparisons with Other Benchmarks
Appendix B Additional Experimental Results
B.1 Fine-Grained Error Analysis
We provide the statistical results on the Code Pass rate under various data categories. The results of Code Pass rate are shown as follows (’Code Pass’ refers to the success rate of code execution).
Table 6: Code Pass rate under various data categories.
Models Prompt Linear w/o Table Linear w/ Table Nonlinear w/o Table Nonlinear w/ Table Overall Pass
Mistral-7B-Instruct-v0.3 zero-shot 5.8 6.3 9.0 0.1 6.9
Qwen2-7b-Instruct zero-shot 21.3 0.1 21.8 12.0 19.2
gpt-4o zero-shot 93.6 93.8 83.5 78.0 90.1
Mistral-7B-Instruct-v0.3 few-shot 87.7 85.0 75.9 84.0 83.8
Qwen2-7b-Instruct few-shot 92.4 85.0 78.9 82.0 87.6
gpt-4o few-shot 94.2 80.0 85.7 86.0 91.7
From the above statistical results, we can observe that the Code Pass rate for nonlinear problems is lower than that for linear problems, indicating that solving nonlinear problems poses certain challenges in terms of coding. Furthermore, we provide 2 case studies to show the errors in code generation by language models, as shown in the following.
## Error Case 1:
import pyscipopt
# Create a new model
model = pyscipopt.Model()
# Define variables
acai_smoothies = model.addVar("acai_smoothies", vtype="I", lb=0)
banana_smoothies = model.addVar("banana_smoothies", vtype="I", lb=0)
# Define objective function
# Minimize the total amount of water used
obj = model.addVar(’obj’)
model.setObjective(obj, "minimize")
model.addCons(obj == 3 * acai_smoothies + 4 * banana_smoothies)
# Add constraints
# Constraint for acai berries
model.addCons(7 * acai_smoothies <= 3500)
# Constraint for banana chocolate
model.addCons(6 * banana_smoothies <= 3200)
# Constraint for more banana smoothies than acai smoothies
model.addCons(banana_smoothies > acai_smoothies)
# Constraint for at least 35% of smoothies being acai berry smoothies
model.addCons(acai_smoothies >= 0.35 * (acai_smoothies + banana_smoothies))
# Solve the problem
model.optimize()
## Error Analysis 1: PySCIPOpt does not support the symbols $<$,$>$; it only supports $\leq$, $\geq$.
## Error Case 2:
import math
import pyscipopt
# Create a new model
model = pyscipopt.Model()
# Define variables for the rescue teams location (x, y)
x = model.addVar(’x’, lb=0) # x must be non-negative
y = model.addVar(’y’)
# Define the mans location
man_x = 3/4
man_y = 0
# Define the objective function: minimize the distance to the man
distance = model.addVar(’distance’)
model.setObjective(distance, "minimize") # We want to minimize the distance
# Add the distance constraint
model.addCons(distance == math.sqrt((x - man_x)**2 + (y - man_y)**2))
# Add the constraint for the swamp: y >= x^2
model.addCons(y >= x**2)
# Solve the problem
model.optimize()
## Error Analysis 2: In PySCIPOpt, you cannot use math.sqrt() directly.
Moreover, we measure the accuracy of the mathematical formulations using the LLM-as-a-judge approach and demonstrate the accuracy of transferring the mathematical formulations to code. We consider this analysis can provide a more comprehensive understanding of the models’ performance. In this experiment, we use deepseek-v2.5 to judge the results.
Table 7: Accuracy of the Mathematical Formulations.
Models Prompt Linear w/o Table Linear w/ Table Nonlinear w/o Table Nonlinear w/ Table Overall Pass
Mistral-7B-Instruct-v0.3 zero-shot 56.4 25.0 23.3 24.0 42.3
Mistral-7B-Instruct-v0.3 few-shot 70.7 38.8 32.3 46.0 56.0
Qwen2-7b-Instruct zero-shot 71.1 40.0 39.1 26.0 56.2
Qwen2-7b-Instruct few-shot 78.7 43.8 41.4 46.0 63.1
deepseek-v2.5 zero-shot 91.8 77.5 63.2 56.0 80.7
deepseek-v2.5 few-shot 93.6 81.3 70.7 68.0 84.8
gpt-4o-mini zero-shot 89.5 72.5 69.2 60.0 80.3
gpt-4o-mini few-shot 90.1 67.5 72.9 70.0 81.7
gpt-4o zero-shot 90.6 82.5 80.5 74.0 85.9
gpt-4o few-shot 92.9 73.8 82.0 70.0 86.1
As shown in the table, small open-source models like Mistral-7B-Instruct-v0.3 and Qwen2-7b-Instruct tend to make more mistakes in mathematical formulations than large models like deepseek-v2.5 and gpt-4o. Moreover, the few-shot setting can significantly decrease the errors of mathematical formulations for all models than the zero-shot setting. These phenomena are also observed in benchmarks like GSM8K and MATH, leading us to conclude that they stem from the model’s capacity for mathematical reasoning.
Table 8: Accuracy of Transferring the Mathematical Formulations to Code.
Models Prompt Linear w/o Table Linear w/ Table Nonlinear w/o Table Nonlinear w/ Table Overall Pass
Mistral-7B-Instruct-v0.3 zero-shot 1.0 0.0 0.0 0.0 0.7
Mistral-7B-Instruct-v0.3 few-shot 50.8 61.3 41.9 39.1 49.9
Qwen2-7b-Instruct zero-shot 4.9 0.0 7.7 0.0 4.7
Qwen2-7b-Instruct few-shot 83.3 62.9 45.5 30.4 72.8
deepseek-v2.5 zero-shot 85.4 87.1 52.4 42.9 77.5
deepseek-v2.5 few-shot 85.0 87.7 57.4 70.6 79.3
gpt-4o-mini zero-shot 85.0 67.2 51.1 56.7 74.4
gpt-4o-mini few-shot 82.8 77.8 19.6 48.6 67.4
gpt-4o zero-shot 86.1 78.8 56.0 57.0 76.9
gpt-4o few-shot 87.1 86.4 61.5 71.4 80.6
For models gpt-4o, gpt-4o-mini, and deepseek-v2.5, their errors in mathematical formulations are relatively few. gpt-4o achieves the sota performance on both two processes. Moreover, as can be seen from Table 8, the code transfer accuracy of Mistral-7B-Instruct-v0.3 and Qwen2-7b-Instruct is significantly lower in the zero-shot setting compared to the few-shot setting, which echoes our previous observation (models with weaker coding abilities are unable to correctly write code relying on their capabilities in a zero-shot setting but can solve problems correctly when provided with few-shot examples).
In addition, we have conducted human-LLM consistency statistics. We randomly select 20 samples and ask 2 master’s degree students in computer science to label them, comparing the labels to those of LLM-as-a-judge. The final results show that 19 out of 20 samples were consistent. This indicates the effectiveness of LLM-as-a-judge in fine-grained error analysis. The prompt of LLM-as-a-judge is shown in Appendix E.1.5.
B.2 Impact of Different Prompting Strategies
We investigate whether step-by-step reasoning then generates the code can help improve model performance. We construct a new prompt (first step reason then write code) named ’few-shot (first reason)’. The experimental results are shown in the following:
Table 9: Performance of Other Prompting Strategy.
Models Prompt Overall Acc
GPT-3.5-Turbo few-shot (ori) 56.4
GPT-3.5-Turbo few-shot (first reason) 55.5 (-0.9)
GPT-4 few-shot (ori) 65.5
GPT-4 few-shot (first reason) 63.8 (-1.7)
GPT-4o-mini few-shot (ori) 55.0
GPT-4o-mini few-shot (first reason) 54.9 (-0.1)
GPT-4o few-shot (ori) 69.4
GPT-4o few-shot (first reason) 67.1 (-2.3)
’few-shot (ori)’ is the original few-shot prompt in out paper, and ’few-shot (first reason)’ is provided in the Appendix E.1.3. It can be seen from the results that ”first understanding the reasoning and then writing code” does not significantly improve performance.
B.3 Pass@k Performance
We provide the Pass@k results as follows (we set the temperature as 0.7) as following. The performance of the model improves with the increase in the number of generation attempts.
Table 10: Pass@k Performance.
Models Pass@5 Pass@10 Pass@15 Pass@20 Pass@25 Pass@30
Llama-3-8B-Instruct 40.5% 55.5% 59.5% 61.5% 62.5% 63.3%
Mistral-7B-Instruct-v0.3 47.1% 57.0% 61.3% 63.8% 64.7% 66.1%
Qwen2-7b-Instruct 59.5% 63.5% 65.8% 67.3% 68.3% 68.6%
B.4 Solving Efficiency of the Generated Codes
To fairly compare with code written by humans, for each tested LLM, we only compare the runtime of the LLM’s code with that of human-written code on the samples where the LLM can provide correct answers. The statistical results are as follows (in seconds):
Table 11: Comparison of Solving Efficiency
Models Prompt LLM code runtime Human code runtime
Mistral-7B-Instruct-v0.3 zero-shot 0.160 0.154
Mistral-7B-Instruct-v0.3 few-shot 0.147 0.151
Qwen2-7b-Instruct zero-shot 0.403 0.157
Qwen2-7b-Instruct few-shot 0.212 0.143
deepseek-v2.5 zero-shot 0.259 0.157
deepseek-v2.5 few-shot 0.146 0.135
gpt-4o-mini zero-shot 0.283 0.159
gpt-4o-mini few-shot 0.114 0.145
gpt-4o zero-shot 0.199 0.157
gpt-4o few-shot 0.152 0.158
The aforementioned outcomes indicate that the efficiency of LLM written code is akin to that of human beings. The principal challenges for LLMs in solving optimization problems pertain to the accurate formulation of mathematical formulations and the generation of code that is devoid of errors.
B.5 Potential of LLMs to Extend Problem Complexity
We conducte the following attempts: We provid a sample to DeepSeek-V2.5 and prompt the LLM to expand the number of variables and constraints of the sample. One of the expansion results is shown in the following:
Original Sample (6 variables, 3 constraints):
## Define Variables:
Gandhi Cloth Company is capable of manufacturing three types of clothing: shirts, shorts, and pants. The manufacture of each type of clothing requires Gandhi to rent the appropriate type of machinery. The company needs to determine the optimal number of each type of clothing to manufacture, and the number of each type of machinery to rent.
// {"number of shirts to manufacture": "Shirt", "range": "Shirt >= 0", "type": "integer"}
// {"number of shorts to manufacture": "Shorts", "range": "Shorts >= 0", "type": "integer"}
// {"number of pants to manufacture": "Pants", "range": "Pants >= 0", "type": "integer"}
// {"number of shirt machinery to rent": "Shirt_Machinery", "range": "Shirt_Machinery >= 0", "type": "integer"}
// {"number of shorts machinery to rent": "Shorts_Machinery", "range": "Shorts_Machinery >= 0", "type": "integer"}
// {"number of pants machinery to rent": "Pants_Machinery", "range": "Pants_Machinery >= 0", "type": "integer"}
## Define Objective Function:
The revenue per shirt is \$12, the revenue per shorts is \$8, and the revenue per pants is \$15.
The cost per shirt is \$6, the cost per shorts is \$4, and the cost per pants is \$8.
The rental cost per shirt machinery per week is \$200, the rental cost per shorts machinery per week is \$150, and the rental cost per pants machinery per week is \$100.
The company wants to maximize the weekly profit.
// Total_Revenue = 12*Shirt + 8*Shorts + 15*Pants
// Total_Cost = 6*Shirt + 4*Shorts + 8*Pants + 200*Shirt_Machinery + 150*Shorts_Machinery + 100*Pants_Machinery
// Objective Function: Maximize: Total_Revenue - Total_Cost
## Generate Constraint-1:
The labor hours required per shirt is 3, the labor hours required per shorts is 2, and the labor hours required per pants is 6. Each week, 150 hours of labor are available.
// 3*Shirt + 2*Shorts + 6*Pants <= 150
## Generate Constraint-2:
The cloth required per shirt is 4 sq yd, the cloth required per shorts is 3 sq yd, and the cloth required per pants is 4 sq yd. Each week, 160 sq yd of cloth are available.
// 4*Shirt + 3*Shorts + 4*Pants <= 160
## Generate Constraint-3:
The company needs to rent at least one machinery for each type of clothing.
// Shirt_Machinery >= 1, Shorts_Machinery >= 1, Pants_Machinery >= 1
Expanded Samples (10 variables, 8 constraints):
### Define Variables:
Gandhi Cloth Company is capable of manufacturing five types of clothing: shirts, shorts, pants, jackets, and skirts. The manufacture of each type of clothing requires Gandhi to rent the appropriate type of machinery. The company needs to determine the optimal number of each type of clothing to manufacture, and the number of each type of machinery to rent.
// {"number of shirts to manufacture": "Shirt", "range": "Shirt >= 0", "type": "integer"}
// {"number of shorts to manufacture": "Shorts", "range": "Shorts >= 0", "type": "integer"}
// {"number of pants to manufacture": "Pants", "range": "Pants >= 0", "type": "integer"}
// {"number of jackets to manufacture": "Jackets", "range": "Jackets >= 0", "type": "integer"}
// {"number of skirts to manufacture": "Skirts", "range": "Skirts >= 0", "type": "integer"}
// {"number of shirt machinery to rent": "Shirt_Machinery", "range": "Shirt_Machinery >= 0", "type": "integer"}
// {"number of shorts machinery to rent": "Shorts_Machinery", "range": "Shorts_Machinery >= 0", "type": "integer"}
// {"number of pants machinery to rent": "Pants_Machinery", "range": "Pants_Machinery >= 0", "type": "integer"}
// {"number of jacket machinery to rent": "Jacket_Machinery", "range": "Jacket_Machinery >= 0", "type": "integer"}
// {"number of skirt machinery to rent": "Skirt_Machinery", "range": "Skirt_Machinery >= 0", "type": "integer"}
### Define Objective Function:
The revenue per shirt is \$12, the revenue per shorts is \$8, the revenue per pants is \$15, the revenue per jacket is \$20, and the revenue per skirt is \$10.
The cost per shirt is \$6, the cost per shorts is \$4, the cost per pants is \$8, the cost per jacket is \$10, and the cost per skirt is \$5.
The rental cost per shirt machinery per week is \$200, the rental cost per shorts machinery per week is \$150, the rental cost per pants machinery per week is \$100, the rental cost per jacket machinery per week is \$250, and the rental cost per skirt machinery per week is \$120.
The company wants to maximize the weekly profit.
// Total_Revenue = 12*Shirt + 8*Shorts + 15*Pants + 20*Jackets + 10*Skirts
// Total_Cost = 6*Shirt + 4*Shorts + 8*Pants + 10*Jackets + 5*Skirts + 200*Shirt_Machinery + 150*Shorts_Machinery + 100*Pants_Machinery + 250*Jacket_Machinery + 120*Skirt_Machinery
// Objective Function: Maximize: Total_Revenue - Total_Cost
### Generate Constraint-1:
The labor hours required per shirt is 3, the labor hours required per shorts is 2, the labor hours required per pants is 6, the labor hours required per jacket is 8, and the labor hours required per skirt is 4. Each week, 1500 hours of labor are available.
// 3*Shirt + 2*Shorts + 6*Pants + 8*Jackets + 4*Skirts <= 1500
### Generate Constraint-2:
The cloth required per shirt is 4 sq yd, the cloth required per shorts is 3 sq yd, the cloth required per pants is 4 sq yd, the cloth required per jacket is 5 sq yd, and the cloth required per skirt is 3 sq yd. Each week, 1600 sq yd of cloth are available.
// 4*Shirt + 3*Shorts + 4*Pants + 5*Jackets + 3*Skirts <= 1600
### Generate Constraint-3:
The company needs to rent at least one machinery for each type of clothing.
// Shirt_Machinery >= 1, Shorts_Machinery >= 1, Pants_Machinery >= 1, Jacket_Machinery >= 1, Skirt_Machinery >= 1
### Additional Constraints:
To ensure the problem remains feasible and interesting, we can add additional constraints based on the production capacity of each machinery type.
#### Constraint-4:
Each shirt machinery can produce up to 20 shirts per week.
// Shirt <= 20 * Shirt_Machinery
#### Constraint-5:
Each shorts machinery can produce up to 30 shorts per week.
// Shorts <= 30 * Shorts_Machinery
#### Constraint-6:
Each pants machinery can produce up to 15 pants per week.
// Pants <= 15 * Pants_Machinery
#### Constraint-7:
Each jacket machinery can produce up to 10 jackets per week.
// Jackets <= 10 * Jacket_Machinery
#### Constraint-8:
Each skirt machinery can produce up to 25 skirts per week.
// Skirts <= 25 * Skirt_Machinery
Furthermore, we had 3 master’s degree students in computer science to examine 10 expanded samples. Specifically, the following points were checked:
Variables: Does the expanded variable relate to the optimization objective? Constraints: Are the expanded constraints reasonable? Formulation: Whether the formulation of the expanded mathematical model is correct. The verification results indicate that 9 out of the 10 expanded samples are correct.
We consider this to be the foundation for the model to self-improve its complex problem-solving abilities. In fact, this is also one of the future work plans we have.
Appendix C More Detail of ReSocratic